TheJag18:

(Mortgage Loans MC) A home buyer is debating between two different mortgages for $167,800. The options are: Option A: 20-year fixed rate loan at 7.45% with a total of $323,198.36 paid in principal and interest over the life of the loan Option B: 15-year fixed rate loan at the same rate. How much more total principal and interest will the buyer pay for the 20-year loan versus the 15-year loan? A spreadsheet was used to calculate the correct answer. Your answer may vary slightly depending on the technology used $46,041.05 $44,061.05 $229,893.95 $292,893.95

Phantomdex:

Okay so: Total payment = Loan amount x (rate/12) x (1 + rate/12)^n/1 + rate/12)^n -1 Loan amount = $167,800 Rate = 7.45% per year (or 0.0745 as a decimal) Number of payments (n) = 15 years * 12 months/year = 180

D14610:

The total amount paid in principal and interest for the 20-year loan is given as $323,198.36. To find out how much the buyer will pay in total for the 15-year loan, we first need to calculate the monthly payment for the 15-year loan. The formula for the monthly pay- ment of a loan is: M = P[r(1 + r)"]/[(1 + r)” − 1] - where: M is the monthly payment, P is the principal loan amount, r is the monthly interest rate (annual rate divided by 12), n is the number of payments (months). Substituting the given values into the formula: r = 7.45/100/12 = 0.00620833 n = 15 * 12 = 180 P = $167,800 M = 167800[0.00620833(1+0.00620833)¹80]/[(1 +0.00620833)¹80 - 1] After calculating the above expression, we find that the monthly payment for the 15- year loan is approximately $1532.17. The total amount paid over the life of the 15-year loan is then the monthly payment times the number of payments: Total = M * n = $1532.17 * 180 = $275,590.60 Finally, to find out how much more the buyer will pay for the 20-year loan versus the 15-year loan, subtract the total amount paid for the 15-year loan from the total amount paid for the 20-year loan: Difference = $323, 198.36 – $275, 590.60 So, the buyer will pay approximately $47, 607.76 more for the 20-year loan versus the 15-year loan.

Phantomdex:

Next: Total payment = $167,800 x (0.0745/12) x (1 + 0.0745/12)^180)/(1 + 0.0745)^180 -1

Phantomdex:

So based on the loan differences. (Which are) Option A: Total paid = $323,198.36 Option B: approximately $269,157.31. So, our correct answer would be $54,041.05 but our other closest option is $46,041.05 which is not exact.

TheJag18:

thx man

Phantomdex:

No problemo

TheJag18:

wait. it cant be 54,041.05

TheJag18:

its not an answer choice and neither is yours D14610

Phantomdex:

Re-read the last part and look at the numbers silly

D14610:

@thejag18 wrote:
its not an answer choice and neither is yours D14610
we're rounding it

TheJag18:

oh. but dont we have to get exact?

Phantomdex:

I said $46,041.05 would be the closest, which is one of the answers

TheJag18:

yeah you did

Phantomdex:

It's the closest you have so.

TheJag18:

ok. thx guys for answering